what is two ways you can answer 25-x=5x-3 without useing subsition
Answer by haileytucki(390)

25-x=5x-3
Since 5x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 5x from both sides.
-x+25-5x=-3
Since -x and -5x are like terms, subtract 5x from -x to get -6x.
-6x+25=-3
Since 25 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 25 from both sides.
-6x=-25-3
Subtract 3 from -25 to get -28.
-6x=-28
Divide each term in the equation by -6.
-(6x)/(-6)=-(28)/(-6)
Simplify the left-hand side of the equation by canceling the common factors.
x=-(28)/(-6)
Simplify the right-hand side of the equation by simplifying each term.
x=(14)/(3)

25-x=5x-3
To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side.
25-x-5x+3=0
Add 3 to 25 to get 28.
28-x-5x=0
Since -x and -5x are like terms, subtract 5x from -x to get -6x.
28-6x=0
Multiply each term in the equation by -1.
-6x*-1+28*-1=0*-1
Simplify the left-hand side of the equation by multiplying out all the terms.
6x-28=0*-1
Multiply 0 by -1 to get 0.
6x-28=0
Since -28 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 28 to both sides.
6x=28
Divide each term in the equation by 6.
(6x)/(6)=(28)/(6)
Simplify the left-hand side of the equation by canceling the common factors.
x=(28)/(6)
Simplify the right-hand side of the equation by simplifying each term.
x=(14)/(3)